# A physically modelled selection function for compact binary mergers in the LIGO-Virgo O3 run and beyond

### v1

@@ 6 Discussion
In this work we have introduced a simple but accurate, physically motivated approximation of the detection probability $P_{\mathrm{det}}$ for binary merger signals in LIGO-Virgo-KAGRA search pipelines analyzing real gravitational-wave data, accounting for dependence on binary masses, orbit-aligned spins and luminosity distance. Observation selection bias, which is quantified via this probability, has a very strong influence on the set of events recovered in any GW catalog, and thus must be taken into account in any kind of population analysis. The expected search sensitivity to a signal with given masses and spins depends on many diverse factors, including the distance and other extrinsic parameters, but also the detector sensitivities, data quality, and the configuration of search pipelines. These effects are most often accounted for by resampling (importance sampling) of simulated signals, i.e. injection sets, but this method is inefficient for general forms of population model, and may incur high uncertainties. Our proposed fit gives the probability of detection as a continuous (and differentiable) function over mass, spin and distance (or redshift), and thus makes it possible to do population studies using models of arbitrarily complex form without increased uncertainty in the selection function. The fit may also be used for “forward modelling” to directly predict the number and distribution of observed events given the outputs of astrophysical population synthesis calculations.
@@
Our fit function is motivated by the known dependence of the signal-to-noise ratio, for a binary at fixed distance, on other parameters, mainly the sky location and binary orientation and the chirp mass. We allow for mass- and spin-dependent corrections to the theoretically expected behaviour: these corrections, as well as an overall scaling (related to the detector network sensitivity), are determined via a maximum-likelihood fit to the O3 injection campaign results. A surprisingly small number of correction terms is needed to obtain stable and physically reasonable results, which accurately predict the number and distribution found injections, as seen by CDF comparisons and KS test statistics. The statistical uncertainty in the fitted $P_{\mathrm{det}}$ is small, except for the high (redshifted) mass limit where detected injections are very sparse. A limitation of our procedure is that we are unable to constrain or verify the behaviour of $P_{\mathrm{det}}$ outside the physical range of masses, distances and spins covered by the injection set, thus care is required if applying the fit in extreme cases such as in the high mass limit.
@@
A version of this fit (without including spin effects) has already been applied in an investigation of the evolution of the BBH population over redshift [21]. While we have concentrated on fitting the results of injections analyzed by search pipelines in real O3 data, the fitting method may also be applied to “semianalytic” injections associated with the O1 and O2 runs [65]. There are, however, significant differences relative to O3 in the method used for identifying these injections as detected – mainly, thresholding on optimal (expected) network SNR rather than on the search pipeline significance, i.e. false alarm rate (FAR) or ranking statistic, as well as the use of constant detector PSDs for each run. Although the resulting integrated sensitivity over O1 and O2 is consistent with total merger rate estimates, these method differences may produce a bias in the distance dependence of the “semianalytic” found injections, compared to what would be found by actual searches of O1 and O2 data with injections. Therefore, rather than a detailed functional fit of O1-O2 semianalytic injections, we instead recommend keeping the O3 fitted parameters, but scaling the $d_{\mathrm{mid}}$ values by a constant factor for each of the O1 and O2 runs (i.e. changing only $D_{0}$ between runs), such that the total counts of found injections for O1 and O2, and the numbers of actual BBH detections (3 and 10, respectively) are consistently predicted.
@@
Concerning methodological extensions, we are currently treating extrinsic angular parameters as unknowns to be averaged over: the form of our sigmoid function over distance, Eq. (4), is primarily determined by how the signal amplitude depends on such nuisance parameters. We can expect, though, that the probability of detection for a binary at a given distance and with known sky location and orientation will be related in a simple way to our fitted parametric form via $d_{\mathrm{mid}}$. Including the explicit angular dependence may facilitate studies of any deviations from isotropy in the merging binary distribution (e.g. [66, 67]).
@@ Acknowledgments
Appendix A Appendix: Injection probability density functions
@@
$p(m_{1})\sim m_{1}^{\alpha}\,,\hskip 14.22636ptm_{1,min}\leq m_{1}\leq m_{1,max}\,,$ (20a)
$p(m_{2}|m_{1})\sim m_{2}^{\beta}\,,\hskip 14.22636ptm_{2,min}\leq m_{2}\leq m_{1}\,.$ (20b)
The BBH injection distribution for the GWTC-3 catalog [3] uses $\alpha=-2.53$, $\beta=1$, $m_{1,min}=m_{2,min}=2\,\mathrm{M_{\odot}}$ and $m_{1,max}=m_{2,max}=100\,\mathrm{M_{\odot}}$. By normalizing Eq. (19) to 1, we obtain the explicit mass distribution
@@
in which $H_{0}=67.9\,\mathrm{km/s/Mpc}$, $\Omega_{m}=0.3065$ and $\Omega_{\Lambda}=1-\Omega_{m}=0.6935$ (following [70]). Thus, the luminosity distance PDF will be [^2] We can easily obtain $\text{d}(d_{L})/\text{d}z$ by evaluating the integrand from (23) at $t=z$.
$p(d_{L})=\frac{p(z)}{\text{d}(d_{L})/\text{d}z}=\frac{p(z)\,H_{0}}{(1+z)c}\sqrt{\Omega_{m}(1+t)^{3}+(1-\Omega_{m})}\,.$ (24)

### v2

@@ 6 Discussion
In this work we have introduced a simple but accurate, physically motivated approximation of the detection probability $P_{\mathrm{det}}$ for binary merger signals in LIGO-Virgo-KAGRA search pipelines analyzing real gravitational-wave data, accounting for dependence on binary masses, orbit-aligned spins and luminosity distance. Observation selection bias, which is quantified via this probability, has a very strong influence on the set of events recovered in any GW catalog, and thus must be taken into account in any kind of population analysis. The expected search sensitivity to a signal with given masses and spins depends on many diverse factors, including the distance and other extrinsic parameters, but also the detector sensitivities, data quality, and the configuration of search pipelines. These effects are most often accounted for by resampling (importance sampling) of simulated signals, i.e. injection sets, but this method is inefficient for general forms of population model, and may incur high uncertainties. Our proposed fit gives the probability of detection as a continuous (and differentiable) function over mass, spin and distance (or redshift), and thus makes it possible to do population studies using models of arbitrarily complex form without increased uncertainty in the selection function [47]. The fit may also be used for “forward modelling” to directly predict the number and distribution of observed events given the outputs of astrophysical population synthesis calculations.
@@
Our fit function is motivated by the known dependence of the signal-to-noise ratio, for a binary at fixed distance, on other parameters, mainly the sky location and binary orientation and the chirp mass. We allow for mass- and spin-dependent corrections to the theoretically expected behaviour: these corrections, as well as an overall scaling (related to the detector network sensitivity), are determined via a maximum-likelihood fit to the O3 injection campaign results. A surprisingly small number of correction terms is needed to obtain physically reasonable results which accurately predict of the number and distribution found injections, as verified by CDF comparisons and KS tests. The statistical uncertainty in the fitted $P_{\mathrm{det}}$ is small, except for the high (redshifted) mass and unequal mass limits where detected injections are very sparse. A limitation of our procedure (common to any empirical estimate of $P_{\mathrm{det}}$ from search results) is that we are unable to constrain or verify the behaviour of $P_{\mathrm{det}}$ outside the physical range of masses, distances and spins covered by the injection set, thus care is required if applying the fit in extreme cases such as the high mass limit. The limitation would of course be lifted by use of a wider injection parameter range.
@@
A version of this fit (without including spin effects) was an essential ingredient for a nonparametric analysis of the evolution of the BBH population over redshift [21]. While we have concentrated on fitting the results of injections analyzed by search pipelines in real O3 data, the fitting method may also be applied to “semianalytic” injections associated with the O1 and O2 runs [81]. There are, however, significant differences relative to O3 in the method used for identifying these injections as detected – mainly, thresholding on optimal (expected) network SNR rather than on the search pipeline significance, i.e. false alarm rate (FAR) or ranking statistic, as well as the use of constant detector PSDs for each run. Although the resulting integrated sensitivity over O1 and O2 is consistent with total merger rate estimates, these method differences may produce a bias in the distance dependence of the “semianalytic” found injections, compared to what would be found by actual searches of O1 and O2 data with injections. Therefore, rather than a detailed functional fit of O1-O2 semianalytic injections, we instead recommend keeping the O3 fitted parameters, but scaling the $d_{\mathrm{mid}}$ values by a constant factor for each of the O1 and O2 runs (i.e. changing only $D_{0}$ between runs), such that the total counts of found injections for O1 and O2, and the numbers of actual BBH detections (3 and 10, respectively) are consistently predicted.
@@
Concerning methodological extensions, ways of marginalizing over or otherwise including uncertainties in $P_{\mathrm{det}}$ are to be investigated: for instance, when performing Bayesian analysis the fit parameters could be sampled over using the associated likelihood. [^6] We thank Colm Talbot for this suggestion. We are currently treating extrinsic angular parameters as unknowns to be averaged over: the form of our sigmoid function over distance in Eq. (4) is primarily determined by how the signal amplitude depends on such nuisance parameters. We expect that the probability of detection for a binary at a given distance and with known sky location and orientation will be related in a simple way to our fitted parametric form via $d_{\mathrm{mid}}$. Including the explicit angular dependence may facilitate studies of any deviations from isotropy in the merging binary distribution (e.g. [82, 83]). Analogously, one could consider quantifying the dependence of $P_{\mathrm{det}}$ on the state of the detector network (rather than simply averaging over time), cf. [64], which may be of interest for detector characterization.
@@ Acknowledgments
Data availability
The data that support the findings of this study are openly available at the following repository: https://github.com/AnaLorenzoMedina/cbc_pdet .
Appendix A Injection probability density functions
@@
$p(m_{1})\sim m_{1}^{\alpha}\,,\hskip 14.22636ptm_{1,min}\leq m_{1}\leq m_{1,max}\,,$ (23)
$p(m_{2}|m_{1})\sim m_{2}^{\beta}\,,\hskip 14.22636ptm_{2,min}\leq m_{2}\leq m_{1}\,.$ (24)
The BBH injection distribution for the GWTC-3 catalog [3] uses $\alpha=-2.53$, $\beta=1$, $m_{1,min}=m_{2,min}=2\,\mathrm{M_{\odot}}$ and $m_{1,max}=m_{2,max}=100\,\mathrm{M_{\odot}}$. By normalizing (22) to 1, we obtain the explicit mass distribution
@@
in which $H_{0}=67.9\,\mathrm{km/s/Mpc}$, $\Omega_{m}=0.3065$ and $\Omega_{\Lambda}=1-\Omega_{m}=0.6935$ (following [86]). Thus, the luminosity distance PDF will be [^7] We can easily obtain $\mathrm{d}(d_{L})/\mathrm{d}z$ by evaluating the integrand from (27) at $t=z$.
$p(d_{L})=\frac{p(z)}{\mathrm{d}(d_{L})/\mathrm{d}z}=\frac{p(z)\,H_{0}}{(1+z)c}\sqrt{\Omega_{m}(1+t)^{3}+(1-\Omega_{m})}\,.$ (28)